OpenStudy (anonymous):

If I Was To Roll A Six Sided Dice , What Is The Possibility It Would Land On 3 ? ( I Need It In A Percent Form , Please ! :P ) @ashneedhelp

OpenStudy (anonymous):

6 possibilities. Rolling a 3 is 1 out of those 6 possibilities. Therefore: 1/6 is the probability. In percent form: 1/6 = x/100 --> x = 16.67 The probability is 16.67%

OpenStudy (anonymous):

Thanks ! :)